# Transformer Neural Opeartors

### 19 Mar 2024 v1

@@ 5 Experiments
Dataset Description and Generation. Two datasets are generated using the TurtleFSI package, one for fluid-structure interaction and the other for fluid dynamics (bergersen2020turtlefsi). We simulate the fluid-structure interaction and the fluid dynamics test cases described above up to time $T=10$, using a constant time-step $\delta t=\frac{T}{n}$, where $n=1000$. The data sets are composed of solution trajectories $[u_{t},p_{t},d_{t}]$ (resp. $[u_{t},p_{t}]$), which denote the approximate solution of the fluid-structure interaction problem (resp. the fluid dynamics problem) at times $t=i\delta t,i\in\{0,\dots,n\}$. These trajectories are obtained on the basis of all triplets $(\mu,c_{1},c_{2})$ describing combinations of fluid viscosities and inlet conditions, such that $\mu\in\{1,5,10\},(c_{1},c_{2})\in\mathcal{I}.$ Additional details regarding the solver and simulation parameters used can be found in the Appendix (see Sec. LABEL:app:dataset).
@@
For both datasets, the pre-trained model is fine-tuned for unseen viscosity $\mu=5.0$ with different numbers of few shot examples. The inlet conditions of these simulations are excluded from the pre-training data. So both viscosity and inlet conditions of the target PDEs are absent in the per-taining dataset. We finetune both of the pre-train models ${\tt\mathcal{G}^{p}_{NS-EW}}$ and ${\tt\mathcal{G}^{p}_{NS}}$ on each dataset.
@@
Figure 5: Test Error of different models for different number of training samples.
@@
Next, when CoDA-NO is pre-trained solely on the Navier-Stokes equation (NS dataset), it shows an impressive ability to adapt to the more challenging fluid-structure interaction PDE (NS+EW dataset). Finally, when CoDA-NO is pre-trained on the more intricate NS+EW dataset, it easily adapts to the simpler NS dataset through fine-tuning.
@@
Interestingly, when more training data is available, CoDA-NO trained from scratch significantly outperforms all other models, including the regular attention mechanism (ViT), U-Net, and GNN. This illustrates the applicability of the CoDA-NO architecture not just for few-shot learning but also in scenarios where abundant data is available, thanks to its effective attention mechanism for function spaces. We also provide an ablation study where we freeze the parameters of the "Encoder" and only train the parameters of the "Predictor". Here we also perform significantly better than other models on most of the dataset (see Appendix Sec. LABEL:app:ablation_forzen).

### 05 Apr 2024 v2

@@ 5 Experiments
Dataset Description and Generation. Two datasets are generated using the TurtleFSI package, one for fluid-structure interaction and the other for fluid dynamics (bergersen2020turtlefsi). We simulate the fluid-structure interaction and the fluid dynamics test cases described above up to time $T=10$, using a constant time-step $\delta t=\frac{T}{n}$, where $n=1000$. The data sets are composed of solution trajectories $[u_{t},p_{t},d_{t}]$ (resp. $[u_{t},p_{t}]$), which denote the approximate solution of the fluid-structure interaction problem (resp. the fluid dynamics problem) at times $t=i\delta t,i\in\{0,\dots,n\}$. These trajectories are obtained on the basis of all triplets $(\mu,c_{1},c_{2})$ describing combinations of fluid viscosities and inlet conditions, such that $\mu\in\{0.5,1,5,10\},(c_{1},c_{2})\in\mathcal{I}.$ashiqAdditional details regarding the solver and simulation parameters used can be found in the Appendix (see Sec. LABEL:app:dataset), where we also discuss the challenges involved and justification of such experiment setup.
@@
[…]
@@
Figure 5: Test Error of different models for different number of training samples. The error along the y-axis is shown in the logarithmic scale.
@@
[…]
@@
Interestingly, when more training data is available, CoDA-NO trained from scratch significantly outperforms all other models, including the regular attention mechanism (ViT), U-Net, and GNN. This illustrates the applicability of the CoDA-NO architecture not just for few-shot learning but also in scenarios where abundant data is available, thanks to its effective attention mechanism for function spaces.
[…]
To showcase the efficiency of all the proposed model components, we have presented a comprehensive analysis of each of them (CoDA-NO layer, VSPE, and Normalization) in Table 3. The results show that the removal of any of these designed components significantly impacts the model’s performance.
[…]
Pretrain
Dataset
# Few Shot Training Samples
5 25 100
NS NS+EW NS NS+EW NS NS+EW
Reg. Atten. - 0.271 0.211 0.061 0.113 0.017 0.020
CoDA-NO - 0.182 0.051 0.008 0.084 0.006 0.004
NO VSPE NS 0.049 0.079 0.009 0.0132 0.004 0.009
NO VSPE NS EW 0.045 0.057 0.010 0.011 0.008 0.004
NO Norm. NS * * 0.023 * 0.008 0.006
NO Norm. NS EW 0.057 0.232 0.012 0.052 0.006 0.006
CoDA-No NS 0.025 0.071 0.007 0.008 0.004 0.005
CoDA-No NS EW 0.024 0.040 0.006 0.005 0.005 0.003
We also provide an ablation study where we freeze the parameters of the "Encoder" and only train the parameters of the "Predictor". Here we also perform significantly better than other models on most of the dataset (see Appendix Sec. LABEL:app:ablation_forzen).
[…]
Adaptation to More Turbulan Flows. We also test the adaptation capability of our model on a more turbulent fluid-solid interaction scenario with viscosity $\mu=0.5$ with Reynold’s number of $4000$. From  Tab. 4, we can observe that, even though the model is pre-trained on data with lower Reynold’s number ($200-2000$), it can seamlessly adapt to more turbulent flow and outperform baselines, which is trained from scratch.
[…]
Pretrain
Dataset
# Few Shot Samples
5 25 100
GINO 0.717 0.292 0.136
DeepO 0.889 0.545 0.259
GNN 0.374 0.310 0.132
ViT 0.878 0.409 0.164
U-net 3.256 0.563 0.292
Ours 0.326 0.264 0.070
Ours NS 0.366 0.161 0.079
Ours NS-EW 0.308 0.143 0.069
[…]
Parameter Count and Computational Cost. Now the present the number of parameters and training/interference time taken by the proposed model along with different baselines used in the study in  Tab. 5. It might seem that models are not compared fairly, as the CoDA-NO has a higher parameter count. However, here, we test the models on a few shot learning problems. Increasing the baselines’ parameter count worsens the overfitting problem.
[…]
To demonstrate this fact, we perform experiments on a fluid-solid interaction dataset with an increased parameter count. We will observe that increasing the parameter count almost always negatively impacts the performance, especially for very few hot learning scenarios (see  Tab. 6). We also note that the additional model parameters and computation are required to learn rich inter-variable dependencies during pre-training and generalize from single to multi-physics during finetuning. Furthermore, the zero-shot super-resolution capability of CoDA-NO is discussed in  Sec. LABEL:app:super_res_test. CoDA-NO is a justified choice due to its seamless adaptation to various PDEs, remarkable performance gap, and zero-shot super-resolution capability, despite having a little more computational overhead.
[…]
Models GNN GINO DeepO ViT Unet CoDA-NO
Inference Time 0.012 0.012 0.006 0.071 0.024 0.440
Training Time 0.136 0.136 0.131 0.273 0.268 1.250
# Parameter ×$\times 10^{6}$ 0.6 60 6 27 30 43
[…]
(Used/High) (Used / High) (Used/ High) (Used / High)
GINO 60/200 0.122 / 0.342 0.053 / 0.066 0.043 / 0.036
DeepO 6 / 25 0.482 / 0.495 0.198 / 0.303 0.107 / 0.083
GNN 0.6/7 0.045 / 0.268 0.009 / 0.031 0.009 / 0.061
ViT 27/100 0.211 / 0.266 0.113 / 0.125 0.020 / 0.022
U-net 30/48 3.579 / 9.462 0.842 / 3.957 0.203 / 0.412